OpenStudy (anonymous):

Which of the following is the solution to this system? 7x - 2y = -3, 14x + y = 14

OpenStudy (anonymous):

Do u have choices?

OpenStudy (anonymous):

\[(4, \frac{ 5 }{ 7 })\] \[(4, \frac{ 7 }{ 5 })\] \[(\frac{ 5 }{ 7 }, 4)\] \[(\frac{ 7 }{ 5 }, 4)\]

OpenStudy (anonymous):

@King_Trae_OKC

OpenStudy (anonymous):

Do u juss want the answer?

OpenStudy (anonymous):

I wanna know how to do it. I think it's just a simple "Plug in & find", but I'm not completely certain.

OpenStudy (anonymous):

Thats exactly what this is

OpenStudy (anonymous):

Alright, thanks. i thought so, it's just been a while since I've done this.

OpenStudy (anonymous):

\(\Huge\sf\color{#ff1ff1}{lol~}{Your~Welcome :)}\)